# Nested Skills in Labor Ecosystems: A Hidden Dimension of Human Capital

### v3

@@ Introduction
By examining three different datasets (median occupational ages, synthetic birth cohorts of individuals, and 70 million job transitions in resumes), we uncover that nested branches are evidence of specialization and career advancement. That is, as individuals progress up the career ladder, they need to acquire and apply skills on nested specialization branches (Fig. 4). Moreover, we find most of the wage premiums for these nested specializations are conditional on foundational, general prerequisite skills they are nested in, unlike unnested specializations without prerequisite skills (Fig. LABEL:fig:wage_curves). This pattern suggests deeply rooted structural disparities in race/ethnicity and gender (Fig. 6). Finally, we examine structural changes in the skill network over time and find a wider gap between nested and unnested branches, suggesting potential barriers to upward mobility (Fig. 7).
@@ Results
These classifications, detailed in SI-Table LABEL:tab:skill_groups, align with our common understanding of general and specialized skill categories. Nevertheless, we ensure the robustness of our findings by testing our results against different group sizes and clustering algorithms (see SI Sec. LABEL:supsec:skill_clustering). In addition to the distribution-based approach, skill generality can also be measured by the median skill levels required across occupations. For example, the median level for general skills is 3.34, for intermediate skills, it is 2.37, and for specific skills, it is 0.87, reflecting the skewed shape of niche skills. In the following, we additionally show that these generality measures are consistent with network-based measures of generality [Mones2012]. Throughout the paper, our results are color-coded for consistency: general (red), intermediary (gray), and specific skills (blue).
@@
Figure 2 (a) shows the backbone of the resulting hierarchical network obtained by aggregating the empirically derived dependencies across all skill pairs. The network extends from general to specialized skills, incorporating their directional dependencies (the full network is shown in Fig. 2 b). Nodes are colored by generality group as in Fig. 1 and positioned based on educational requirements (x-axis) and Local Reaching Centrality (y-axis), a measure of skill generality denoting the number of other skills reachable from the focal skill [Mones2012]. The network reveals distinct specialization paths and a partially nested architecture. Methods and SI Sec. LABEL:supsec:conditional_dependencies provide detailed parameters for statistical filterings and the threshold for directionality and backbone structure for Fig. 2.
Constructing a network structure from these conditional directions provides a methodologically consistent definition of general and specific skills using reaching centrality [Mones2012] as an alternative measure for generality, as this can reflect the mass of interdependent nodes on the focal node (0.71 correlated). Chains of dependencies for select examples are also well embedded as expected, such as Deductive Reasoning to Math skills to Programming, exemplifying the nesting of skills in the skill hierarchy. Negotiation has a different set of dependencies compared to Programming, including Systems Analysis. Supplementary Information Secs. LABEL:supsec:RN_vs._NP and LABEL:supsec:hispanic_skill_entrapment offer brief case studies highlighting the role of dependency chains in career progress and specialization. Finally, we include the fully labeled visualizations of Fig. 2 (a-b) in SI Figs. LABEL:fig:figure_2b_labeled and LABEL:fig:full_figure_2b_labeled for further examinations.
@@
To systematically quantify and differentiate these observations on skills, we introduce the ecological measure of nestedness and individual contribution scores where specialist species engage preferentially with generalists [Bascompte2003]. This analogy extends to the skill ecosystem, where general human capital forms the bedrock for the acquisition and application of more specialized skills [Saavedra2011, Saavedra2009]. Therefore, we first measure an overarching nested structure in human capital $N$. There are a number of different ways to measure nested structures. We employ several measures commonly used in ecology, such as the overlap index ($N_{c}$), checkerboard score, Temperature, and NODF, to ensure the analysis withstands the test of different nestedness measurements [Stone1990, Almeida-neto2008, write1992, write1992, Saavedra2011] (See SI Sec. LABEL:suppsec:nestedness for the full analyses and robustness tests).
@@
The advantage of the second dataset is the information on both the age and demographics of individuals, allowing us to decompose our findings by gender. Differentiating skill trends by gender uncovers a gap in specializations that emerges around 30. Men continue to grow their general and nested skills until their 50s, whereas for women, the increase in these skills hits a plateau in their early 30s, the typical age range for first-time mothers in the US. Supplementary Information Secs. LABEL:supsec:Parenthood_Male_vs_Female and LABEL:supsec:female_job_sorting further investigate the influence of parenthood on male and female workers by slicing data by those with and without children as well as the influence of sorting into jobs based on schedule and working hours, respectively. These findings are robust to conditioning out yearly economic conditions (SI Fig. LABEL:fig:individuals'_age_and_skill_-_year_effects). In the following sections and in Fig. LABEL:fig:Skill_Age_Gender_Race_Trends_-_year_effects, we offer more detailed breakdowns of these gender disparity trends with respect to race and ethnicity. Notably, education does not fully account for the growth in skill documented by our analysis. As SI Fig. LABEL:fig:individuals'_age_and_skill_and_education shows, the share of educational attendance is negligible after the age of 30, while skill growth continues. Similar patterns, in more modest magnitudes, emerge for workers with no more than high school diplomas (SI Fig. LABEL:fig:individuals'_age_and_skill_-_no_college.)
@@
Figures 4 (g-h) show the average skill levels required in job sequences held across career paths, and Fig. 4 (i) displays changes in skill requirements for the $i$th job transition, $\Delta_{i}$, excluding job transitions within the same occupation ($\Delta_{i}=0$). On aggregate, career journeys unfold with increasing stocks of both general and nested skills ($c_{s}>0$), suggesting that nested specialization paths require simultaneous increases in nested specific skills along with their dependency skills. In addition, we find skill portfolios typically stabilize within the first five job transitions ($\Delta_{i>5}\approx 0$), and in the first three jobs ($i<3$), nested skills require more general skills than later ($\Delta^{general}_{i<3}\gg\Delta^{nested}_{i<3}$), after which they become comparable ($\Delta^{general}_{i>3}\approx\Delta^{nested}_{i>3}$). The continued growth in general skills across career paths suggests that these skills need to be continuously enhanced regardless of career stage. As a benchmark, we create bootstrapped job sequences (gray marks around zero) that randomize the order of jobs as if there were no career development, confirming that the observed trends are indeed attributed to career developments (see SI Sec. LABEL:supsec:_bootstrapping_BG for details).
@@
To explore nested specialization, we choose registered nurses (RNs) and nurse practitioners (NPs) by analyzing resume data to understand how skill and wage differences manifest in career trajectories. Supplementary Information Fig. LABEL:fig:RN_vs._NP shows the additional skills (necessary to prescribe medicine and diagnostic tests) in higher-paying NP positions appear in nested paths with growth in both general and dependent niche skills such as medicine, therapy, biology, science, and chemistry (see SI-Sec. LABEL:supsec:RN_vs._NP for the detailed analysis). In addition, SI Sec. LABEL:supsec:hispanic_skill_entrapment makes a case wherein insufficient levels of certain general skills preclude the development of the dependent niche skills, once again highlighting how our framework teases out pathways for developing human capital.
@@
All three empirical observations consistently depict nested specializations (i.e., growth in both general and nested skills) throughout career trajectories, while unnested skills are left relatively underdeveloped. The resume analysis offers direct evidence of a recurring yet counterintuitive pattern: valuable specialization is not just about developing niche skills; it is conditional on advancing the required more general skills. This suggests that the conventional model, where basic general skills precede advanced specialized skills, is not entirely accurate. Instead, career paths tend to unfold with increasing emphasis on general skills and their dependent, nested skills. While research has emphasized the role of education, Fig. 4 (and SI Figs. LABEL:fig:individuals'_age_and_skill_and_education and LABEL:fig:individuals'_age_and_skill_-_no_college) reveal that skill advancement continues long after the age of schooling, suggesting nested specialization pathways operate through but also beyond education [Jovanovic1997, Mincer1974, Arrow1962, Lucas1988, Hermo2022], challenging the commonly held role of education in developing human capital.
@@
One might argue that our findings are driven by management/administration jobs, which are typically undertaken later in careers with higher wages. To ensure they do not drive our findings, we repeated the entire analysis without these factors and found consistent results (see SI Sec. LABEL:sec:robustness_check:_no_managers). Also, we repeated the entire analysis, excluding social skills, and again our results remained robust, suggesting that our findings are attributed to the intrinsic structure of skills rather than the influence of particular social skills or managerial jobs (see SI Sec. LABEL:sec:social_skills for the full analyses).
@@
Further analyses in SI Sec. LABEL:supsec:_add_-_returns_to_skill demonstrate that these results are robust to controlling for education, training, and workplace experience and hold across subsamples of major occupational groups. Again, the results are not driven by managerial occupations or social skills, usual suspect factors in wage premium (see the results in SI-Table LABEL:tab:wage_reg_on_skill_endowment, and SI Figs. LABEL:fig:SI_education_skill_level-LABEL:fig:SI_wage_skill_level, LABEL:fig:Figure_3_full_|_major_occupation_groups, LABEL:fig:returns_to_skills_hierachy_gen_dependence_cor_no_manager, and LABEL:fig:social_skills).
@@
Figure 6: Skill Disparity in Demographic Distribution of race/ethnicity and gender (a) The relative average skill level, education level, and weekly wages for Asian, Black, and Hispanic/Latinx workers compared to White workers (expressed as a ratio). (b) The relative average skill level, education level, and weekly wages for female workers compared to male workers. 95% confidence intervals for each estimated ratio are calculated by bootstrapping subsamples (see Methods). These differentials are robust to measurement (SI Fig. LABEL:fig:Tong_et_al_race_gender_skill_distribution), follow similar age trends seen in Fig. 4, and are robust to time-variant economic factors (Fig. LABEL:fig:Skill_Age_Gender_Race_Trends_-_year_effects.) SI Figs. LABEL:fig:Temporal_Race_Gaps_-_Skills,_Education,_Wages and LABEL:fig:Temporal_Gender_Gaps_-_Skills,_Education,_Wages, further show the gaps have narrowed over time.
@@
We explore this further in a brief case study of how language-skill requirements may keep workers out of jobs that require certain nested skills (see Supplementary Information Sec. LABEL:supsec:hispanic_skill_entrapment). To do so, we leverage the hierarchical nature of our skill network. This allows us to distinguish between nested skills that depend on (general) language skills and nested skills that don’t. Within the group of Hispanic workers, we find particularly large gaps in language-dependent nested skills compared to other nested skills for workers who have recently moved to the US. Such workers may instead develop un-nested skills, leading to “skill traps” that are associated with long-run wage penalties (SI Fig. LABEL:fig:wage_curves). Taken together, these findings indicate that closing wage gaps for Black workers may require different solutions than for Hispanic workers.
Figure 6 (b) focuses on skill gaps between men and women across social groups. The most pronounced disparities exist in nested and unnested specializations. Except for in the Asian subsample, women tend to work in occupations that require higher levels of education and general skills than men. However, this does not translate into higher levels of nested skills, where women tend to fall behind men. These disparities are likely to contribute to the well-known gender wage gap we observe in the right-most panel. Encouragingly, this gap has narrowed over time, as demonstrated in SI Fig. LABEL:fig:Temporal_Gender_Gaps_-_Skills,_Education,_Wages. However, the disconnect between education and general skills on the one hand and wages and nested skills on the other is puzzling. Supplementary Information Sections LABEL:supsec:Parenthood_Male_vs_Female and LABEL:supsec:female_job_sorting probe deeper into these gender gaps. This analysis suggests that parenthood, as well as the fact that women often work in jobs with more regular and predictable work schedules, impact both wages and skill development [Bertrand2009, Goldin2015, Canon2016]. In fact, whereas having children is associated with reduced general and nested skills for women, men with children tend to have higher levels of general and nested skills than men without children in the household. When it comes to work schedules, Similarly, we find that the association between gender and nested skill requirements at work is reduced by over a third when we control for irregular hours and overtime in an occupation.
Finally, Section LABEL:section:_add_-_geographical_distribution_of_skills of the Supplementary Information studies the geographic distribution of skills, showing that general skills concentrate in densely populated urban areas. This finding is in line with prior work that highlights the diverse and complex economic activities that are found in large urban economies [Youn2016, gomez2016explaining, Hong2020, Balland2020, Bettencourt2014, Gomez-Lievano2021]. Moreover, this greater concentration of general skills in large cities can account for about one-third of the well-established urban wage premium [Glaeser2001].
@@
Figure 7 (a) indicates an increase in the demand for general skills, as evidenced by the shift from the dotted to the solid distribution. This increase in demand for general skills corresponds to higher wage premiums over the recent decade, suggesting that the economy has been rewarding workers with a broad set of skills (see SI Fig. LABEL:fig:Wage_and_education_2003). However, the seemingly static distribution of specific skills masks underlying changes in the application of nested and unnested skills. As shown in Figure 7(b), there has been a rise in the application of nested skills and a decline in the use of unnested skills between 2005 and 2019, reiterating the importance of considering skill interdependencies when analyzing changes in the skill structure.
@@
However, the shift towards greater overall nestedness has been uneven across different skill sets. Figure 7 (c) demonstrates this as a widening gap between the nested and unnested branches over the decades. Supplementary Fig. LABEL:fig:Figure_7_supp (b) shows a strengthening in the connections among nested skills, indicating their growing complexity and mutual dependence, while SI Fig. LABEL:fig:Figure_7_supp (d) reveals weakened dependency chains of unnested skills. All in all, this trend is observed as a broadening and deepening of nested skill branches within the hierarchy in Fig.  7 (c), reflecting an increase in the complexity and interdependence of these specialized skill areas [DemingDavidJ2020EDCJ, Tong2021].
Indeed, the chasm between the two types of specializations has alarmingly broadened within the educational domain over the last two decades [Xu2021, Lin2022, Althobaiti2022]. In response to this, the structural changes in the hierarchical tree network are concerning, given the significance of these specializations for future career developments and wage premiums. They reveal an economy wherein the structure of valuable human capital has grown more nested, reinforcing barriers to workers without the necessary fundamental skills, who are often entrapped in unnested specialization pathways (see SI Figs. LABEL:fig:wage_curves and LABEL:fig:hispanics_language_skills). The widening gap between the nested and unnested specialization paths could indicate strongly rooted chronic disparity.

### v4

@@ Introduction
By examining three different datasets (median occupational ages, synthetic birth cohorts of individuals, and 70 million job transitions in resumes), we uncover that nested branches are evidence of specialization and career advancement. That is, as individuals progress up the career ladder, they need to acquire and apply skills on nested specialization branches (Fig. 4). Moreover, we find most of the wage premiums for these nested specializations are conditional on foundational, general prerequisite skills they are nested in, unlike unnested specializations without prerequisite skills (Fig. S30). This pattern suggests deeply rooted structural disparities in race/ethnicity and gender (Fig. 6). Finally, we examine structural changes in the skill network over time and find a wider gap between nested and unnested branches, suggesting potential barriers to upward mobility (Fig. 7).
@@ Results
These classifications, detailed in SI-Table LABEL:tab:skill_groups, align with our common understanding of general and specialized skill categories. Nevertheless, we ensure the robustness of our findings by testing our results against different group sizes and clustering algorithms (see SI Sec. 1). In addition to the distribution-based approach, skill generality can also be measured by the median skill levels required across occupations. For example, the median level for general skills is 3.34, for intermediate skills, it is 2.37, and for specific skills, it is 0.87, reflecting the skewed shape of niche skills. In the following, we additionally show that these generality measures are consistent with network-based measures of generality [Mones2012]. Throughout the paper, our results are color-coded for consistency: general (red), intermediary (gray), and specific skills (blue).
@@
Figure 2 (a) shows the backbone of the resulting hierarchical network obtained by aggregating the empirically derived dependencies across all skill pairs. The network extends from general to specialized skills, incorporating their directional dependencies (the full network is shown in Fig. 2 b). Nodes are colored by generality group as in Fig. 1 and positioned based on educational requirements (x-axis) and Local Reaching Centrality (y-axis), a measure of skill generality denoting the number of other skills reachable from the focal skill [Mones2012]. The network reveals distinct specialization paths and a partially nested architecture. Methods and SI Sec. 3 provide detailed parameters for statistical filterings and the threshold for directionality and backbone structure for Fig. 2.
Constructing a network structure from these conditional directions provides a methodologically consistent definition of general and specific skills using reaching centrality [Mones2012] as an alternative measure for generality, as this can reflect the mass of interdependent nodes on the focal node (0.71 correlated). Chains of dependencies for select examples are also well embedded as expected, such as Deductive Reasoning to Math skills to Programming, exemplifying the nesting of skills in the skill hierarchy. Negotiation has a different set of dependencies compared to Programming, including Systems Analysis. Supplementary Information Secs. 3.4 and 3.5 offer brief case studies highlighting the role of dependency chains in career progress and specialization. Finally, we include the fully labeled visualizations of Fig. 2 (a-b) in SI Figs. S15 and S14 for further examinations.
@@
To systematically quantify and differentiate these observations on skills, we introduce the ecological measure of nestedness and individual contribution scores where specialist species engage preferentially with generalists [Bascompte2003]. This analogy extends to the skill ecosystem, where general human capital forms the bedrock for the acquisition and application of more specialized skills [Saavedra2011, Saavedra2009]. Therefore, we first measure an overarching nested structure in human capital $N$. There are a number of different ways to measure nested structures. We employ several measures commonly used in ecology, such as the overlap index ($N_{c}$), checkerboard score, Temperature, and NODF, to ensure the analysis withstands the test of different nestedness measurements [Stone1990, Almeida-neto2008, write1992, write1992, Saavedra2011] (See SI Sec. 2 for the full analyses and robustness tests).
@@
The advantage of the second dataset is the information on both the age and demographics of individuals, allowing us to decompose our findings by gender. Differentiating skill trends by gender uncovers a gap in specializations that emerges around 30. Men continue to grow their general and nested skills until their 50s, whereas for women, the increase in these skills hits a plateau in their early 30s, the typical age range for first-time mothers in the US. Supplementary Information Secs. 7.1 and 7.2 further investigate the influence of parenthood on male and female workers by slicing data by those with and without children as well as the influence of sorting into jobs based on schedule and working hours, respectively. These findings are robust to conditioning out yearly economic conditions (SI Fig. S27). In the following sections and in Fig. S53, we offer more detailed breakdowns of these gender disparity trends with respect to race and ethnicity. Notably, education does not fully account for the growth in skill documented by our analysis. As SI Fig. S28 shows, the share of educational attendance is negligible after the age of 30, while skill growth continues. Similar patterns, in more modest magnitudes, emerge for workers with no more than high school diplomas (SI Fig. S29.)
@@
Figures 4 (g-h) show the average skill levels required in job sequences held across career paths, and Fig. 4 (i) displays changes in skill requirements for the $i$th job transition, $\Delta_{i}$, excluding job transitions within the same occupation ($\Delta_{i}=0$). On aggregate, career journeys unfold with increasing stocks of both general and nested skills ($c_{s}>0$), suggesting that nested specialization paths require simultaneous increases in nested specific skills along with their dependency skills. In addition, we find skill portfolios typically stabilize within the first five job transitions ($\Delta_{i>5}\approx 0$), and in the first three jobs ($i<3$), nested skills require more general skills than later ($\Delta^{general}_{i<3}\gg\Delta^{nested}_{i<3}$), after which they become comparable ($\Delta^{general}_{i>3}\approx\Delta^{nested}_{i>3}$). The continued growth in general skills across career paths suggests that these skills need to be continuously enhanced regardless of career stage. As a benchmark, we create bootstrapped job sequences (gray marks around zero) that randomize the order of jobs as if there were no career development, confirming that the observed trends are indeed attributed to career developments (see SI Sec. 4.1 for details).
@@
To explore nested specialization, we choose registered nurses (RNs) and nurse practitioners (NPs) by analyzing resume data to understand how skill and wage differences manifest in career trajectories. Supplementary Information Fig. S17 shows the additional skills (necessary to prescribe medicine and diagnostic tests) in higher-paying NP positions appear in nested paths with growth in both general and dependent niche skills such as medicine, therapy, biology, science, and chemistry (see SI-Sec. 3.4 for the detailed analysis). In addition, SI Sec. 3.5 makes a case wherein insufficient levels of certain general skills preclude the development of the dependent niche skills, once again highlighting how our framework teases out pathways for developing human capital.
@@
All three empirical observations consistently depict nested specializations (i.e., growth in both general and nested skills) throughout career trajectories, while unnested skills are left relatively underdeveloped. The resume analysis offers direct evidence of a recurring yet counterintuitive pattern: valuable specialization is not just about developing niche skills; it is conditional on advancing the required more general skills. This suggests that the conventional model, where basic general skills precede advanced specialized skills, is not entirely accurate. Instead, career paths tend to unfold with increasing emphasis on general skills and their dependent, nested skills. While research has emphasized the role of education, Fig. 4 (and SI Figs. S28 and S29) reveal that skill advancement continues long after the age of schooling, suggesting nested specialization pathways operate through but also beyond education [Jovanovic1997, Mincer1974, Arrow1962, Lucas1988, Hermo2022], challenging the commonly held role of education in developing human capital.
@@
One might argue that our findings are driven by management/administration jobs, which are typically undertaken later in careers with higher wages. To ensure they do not drive our findings, we repeated the entire analysis without these factors and found consistent results (see SI Sec. 9.1). Also, we repeated the entire analysis, excluding social skills, and again our results remained robust, suggesting that our findings are attributed to the intrinsic structure of skills rather than the influence of particular social skills or managerial jobs (see SI Sec. 9.2 for the full analyses).
@@
Further analyses in SI Sec. 5 demonstrate that these results are robust to controlling for education, training, and workplace experience and hold across subsamples of major occupational groups. Again, the results are not driven by managerial occupations or social skills, usual suspect factors in wage premium (see the results in SI-Table S5, and SI Figs. S31-S33, S38, S61, and S63).
@@
Figure 6: Skill Disparity in Demographic Distribution of race/ethnicity and gender (a) The relative average skill level, education level, and weekly wages for Asian, Black, and Hispanic/Latinx workers compared to White workers (expressed as a ratio). (b) The relative average skill level, education level, and weekly wages for female workers compared to male workers. 95% confidence intervals for each estimated ratio are calculated by bootstrapping subsamples (see Methods). These differentials are robust to measurement (SI Fig. S48), follow similar age trends seen in Fig. 4, and are robust to time-variant economic factors (Fig. S53.) SI Figs. S50 and S51, further show the gaps have narrowed over time.
@@
We explore this further in a brief case study of how language-skill requirements may keep workers out of jobs that require certain nested skills (see Supplementary Information Sec. 3.5). To do so, we leverage the hierarchical nature of our skill network. This allows us to distinguish between nested skills that depend on (general) language skills and nested skills that don’t. Within the group of Hispanic workers, we find particularly large gaps in language-dependent nested skills compared to other nested skills for workers who have recently moved to the US. Such workers may instead develop un-nested skills, leading to “skill traps” that are associated with long-run wage penalties (SI Fig. S30). Taken together, these findings indicate that closing wage gaps for Black workers may require different solutions than for Hispanic workers.
Figure 6 (b) focuses on skill gaps between men and women across social groups. The most pronounced disparities exist in nested and unnested specializations. Except for in the Asian subsample, women tend to work in occupations that require higher levels of education and general skills than men. However, this does not translate into higher levels of nested skills, where women tend to fall behind men. These disparities are likely to contribute to the well-known gender wage gap we observe in the right-most panel. Encouragingly, this gap has narrowed over time, as demonstrated in SI Fig. S51. However, the disconnect between education and general skills on the one hand and wages and nested skills on the other is puzzling. Supplementary Information Sections 7.1 and 7.2 probe deeper into these gender gaps. This analysis suggests that parenthood, as well as the fact that women often work in jobs with more regular and predictable work schedules, impact both wages and skill development [Bertrand2009, Goldin2015, Canon2016]. In fact, whereas having children is associated with reduced general and nested skills for women, men with children tend to have higher levels of general and nested skills than men without children in the household. When it comes to work schedules, Similarly, we find that the association between gender and nested skill requirements at work is reduced by over a third when we control for irregular hours and overtime in an occupation.
Finally, Section 6 of the Supplementary Information studies the geographic distribution of skills, showing that general skills concentrate in densely populated urban areas. This finding is in line with prior work that highlights the diverse and complex economic activities that are found in large urban economies [Youn2016, gomez2016explaining, Hong2020, Balland2020, Bettencourt2014, Gomez-Lievano2021]. Moreover, this greater concentration of general skills in large cities can account for about one-third of the well-established urban wage premium [Glaeser2001].
@@
Figure 7 (a) indicates an increase in the demand for general skills, as evidenced by the shift from the dotted to the solid distribution. This increase in demand for general skills corresponds to higher wage premiums over the recent decade, suggesting that the economy has been rewarding workers with a broad set of skills (see SI Fig. S39). However, the seemingly static distribution of specific skills masks underlying changes in the application of nested and unnested skills. As shown in Figure 7(b), there has been a rise in the application of nested skills and a decline in the use of unnested skills between 2005 and 2019, reiterating the importance of considering skill interdependencies when analyzing changes in the skill structure.
@@
However, the shift towards greater overall nestedness has been uneven across different skill sets. Figure 7 (c) demonstrates this as a widening gap between the nested and unnested branches over the decades. Supplementary Fig. S56 (b) shows a strengthening in the connections among nested skills, indicating their growing complexity and mutual dependence, while SI Fig. S56 (d) reveals weakened dependency chains of unnested skills. All in all, this trend is observed as a broadening and deepening of nested skill branches within the hierarchy in Fig.  7 (c), reflecting an increase in the complexity and interdependence of these specialized skill areas [DemingDavidJ2020EDCJ, Tong2021].
Indeed, the chasm between the two types of specializations has alarmingly broadened within the educational domain over the last two decades [Xu2021, Lin2022, Althobaiti2022]. In response to this, the structural changes in the hierarchical tree network are concerning, given the significance of these specializations for future career developments and wage premiums. They reveal an economy wherein the structure of valuable human capital has grown more nested, reinforcing barriers to workers without the necessary fundamental skills, who are often entrapped in unnested specialization pathways (see SI Figs. S30 and S19). The widening gap between the nested and unnested specialization paths could indicate strongly rooted chronic disparity.